This page is used to inform Network EQ customers of service status information. It is hosted on a different network off-site to our main web site to ensure it is always available. You may like to bookmark this page for future reference.

Our main website can be found at www.networkeq.co.uk.

If you need technical assistance please email support@networkeq.co.uk or call us on +44 (0)333 3441270.

We are currently experiencing an issue with some of our virtual servers, initial investigation indicates this is due to a storage issue.

We are working to restore service as quickly as possible and will post further updates here as we have them.

UPDATE 15:30 - There was an issue with the storage volumes on one of our SAN controllers, these have now been put back online and we are restarting affected customer servers. This process may take a while but all servers should be back online soon.

UPDATE 15:52 - All customer servers should now be back online (although some are still in the process of booting up). We will manually check each server to ensure it's working but if you are still having issues after 16:00 please raise a support ticket.

We are currently experiencing an issue with connectivity at our primary data centre resulting in the unavailability of customer servers.

UPDATE 10:52 - We have been informed that there is a power issue at our primary data centre which explains the unscheduled downtime. Engineers at the data centre are looking into this as a matter of urgency and we'll update this page as soon as we have further information.

UPDATE 11:12 - The power has just been restored at the data centre although our equipment hasn't yet come back online so we are continuing to work with our data centre provider to resolve the situation.

UPDATE 11:21 - The data centre provider is working to bring services back online and we have a support ticket open with them to track the progress of this. We will post further details on the issue and the cause in due course.

UPDATE 11:45 - Power has been restored to our equipment but the firewalls haven't yet come back online. We are working with remote hands at the data centre to diagnose the issue and attempt to bring services back online.

UPDATE 11:59 - We are setting up console access to the firewalls in order to diagnose the problem and attempt to bring them back online.

UPDATE 12:14 - We have console access to the firewalls and our network engineers are working to determine why the network is not functioning correctly after the earlier power issue.

UPDATE 12:31 - We have remote access to the firewalls and we are continuing to diagnose the networking issue.

UPDATE 12:41 - We are continuing to determine which parts of our infrastructure are running correctly thereby eliminating possible issues. The virtualisation platform is partially working and some virtual machines are running although still inaccessible.

UPDATE 12:48 - Customer servers are now being restarted and the infrastructure appears to be operating correctly. It may take a while for all machines to come back online and we will work through all servers to ensure they have booted correctly and are working. If you are still having issues after 13:30 please raise a ticket with us and we will investigate further.

UPDATE 13:08 - All customer servers are now back online with the exception of spam filtering which we are working on resolving now. Emails will be queued until this is resolved. Please accept our apologies for the extended downtime caused by the power issue at our data centre. We will investigate the cause and issue a full report in due course. If you are still having issues with your service please raise a support ticket or call us on 0333 3441270.

UPDATE 13:43 - Spam filtering is now working again so all customer services should be restored.

One of our shared hosting servers, cpanel02, required an emergency reboot and is now running a file system check. Due to the size of the file system this could take some time.

We apologise for any inconvenience caused.

We are currently experiencing a disruption to our inbound email services. This is effecting all users of our email service.

All emails sent to our customers using this service will not be received at present. However these emails will not be lost, as they will be queued for re-delivery.

Outbound email is unaffected.

We are working to restore full services as quickly as possible. We apologize for any inconvenience caused.

We are currently experiencing a storage issue that has caused a small number of customer virtual machines to become unavailable. One of the disk groups on our storage network has become full unexpectedly this evening, causing the storage to go offline for some customers.

We are currently working to resolve the issue by bringing the affected disks back online and restarting the affected virtual machines.

We will also be investigating the underlying issue and improving our monitoring in due course to ensure that this situation will not reoccur.

UPDATE 21:30: All affected customer machines are back online and we are investigating the underlying cause as to why the disk group was becoming full.

We experienced an Internet connectivity issue at our primary data centre, which affected most customer services for 20 minutes. We have confirmed this was caused by an issue with our upstream provider's network and are awaiting further information.

The network has now been stable for over 30 minutes and we are not expecting any further disruption. We apologise for any inconvenience caused.

Today around 15:15 PM it was necessary to restart one of our shared hosting servers. This may have caused some unexpected disruption for some of ours users. All service should now be restored and we will investigate as to the cause.

We apologize for any inconvenience caused.

During this Window we will be upgrading the XenSever hypervisor software that powers our virtualisation environment. We will utilise the XenMotion funtionality during this window to move virtual machines between hosts, but the environment is at risk for the duration of the window as we will have reduced redundancy and the upgrade carries a small risk. The upgrade is necessary for us to install security updates and to support new operating systems such as Windows Server 2012 R2.

After the upgrade is complete virtual machines will need to be rebooted to install updated XenServer tools; if this affects you then our support team will contact you separately to schedule this reboot. Shared hosting servers will be rebooted once the upgrade is complete.

We apologise for any inconvenience caused.

During this window we will be performing scheduled maintenance to upgrade the MySQL version on one of our shared Linux hosting servers, cpanel02. During this window websites may be temporarily inaccessible for short periods. The server will also be restarted to install kernel updates.

We apologise for any inconvenience caused.

During this window we will be performing scheduled maintenance to upgrade the software on our firewalls. Connections may be interrupted for up to one minute during this upgrade.

We apologise for any inconvenience caused.